### v2

@@ 2.1 Point cloud completion
The pioneer work of point cloud completion is PCN [52], which proposed a coarse-to-fine approach that is widely referenced in following studies [25, 21, 34, 28, 42, 57, 58, 44, 55, 37, 39, 38, 43, 33, 23, 3]. Though there are differences in feature extraction and utilization, the basic idea of these studies is to reconstruct the skeleton of the complete shape first and then refine it. PointTr [50] does not follow the coarse-to-fine manner but only generates the missing part of the partial point cloud. The idea of only generating the missing part also appeared in [49, 51]. In [32], a generative adversarial network is used for point cloud completion. PMP-Net [40] and PMP-Net++ [41] treat point cloud completion as a kind of deformation and complete point cloud by moving points to the right positions. P2C [4] introduces additional losses to supervise the latent expressions. Other unsupervised point cloud completion works [12, 53, 17] achieve higher robustness through special training strategies Limited by inputs, these models must learn information about the complete shape from the occluded shape, which may lead to turning the task into a translation process from the occluded shape to the complete shape without meticulous analysis and design of the model.
@@ 4 Experimental Results
In this section, we first introduce the dataset and evaluation metrics in section 4.1. Quantitative and Qualitative comparisons are shown in section 4.2. Ablation studies are conducted in section 4.3. We also report the generalization ability of our method in section 4.4. Finally, the model complexity is shown in section 4.5.
@@ 4.4.1 Results on Unknown categories of ShapeNet-ViPC
Methods Number of parameters(M) $CD\times 10^{3}$
ViPC[56] 17.43 3.308
CSDN[59] 16.85 2.570
XMFnet[1] 9.57 1.443
Ours 9.03 1.211
Table 5: Comparisons on model sizes and performance.
@@ 4.4.2 Results on Real Scenes
We also report qualitative results on KITTI [11] cars extracted by [52]. Qualitative results are shown in Fig. 8. Though there is a synthetic-to-real gap, our method can still able to give a reasonable prediction.
4.5 Comparisons about Model Sizes
We compare the size of our model with existing view-guided completion models[56, 59, 1] in table 5. The comparison results show that our model achieves the best results (+ 16% CD) with the smallest number of parameters.
@@ ACKNOWLEDGMENT
A. Implementation Details
Input partial point cloud contains 2048 points, and input single view image is a $224\times 224\times 3$ RGB image. The output complete point cloud contains 2048 points and is evaluated with ground truth point cloud of 2048 points.
The train-test split follows the list provided by ViPC [56]. For each object, ShapeNet-ViPC [56] provides 24 different views corresponding to different missing situations. During training and testing, pairs of point clouds and images are randomly fed to the model. This means that a partial point cloud may be assisted by different images in different training epochs.
We use the Adam optimizer with an initial learning rate of 0.001 to train our model. We train our model for 160 epochs. The learning rate decreased by 70% every 16 epochs.
During tokenization, the point cloud is downsampled to 256 points and image is divided by a $16\times 16$ grid. Therefore the number of point cloud tokens and image tokens are both 256. Each token in the token feature has 192 channels. The token features keep the size of $256\times 192$ when passing the SFE and SFTnet. The fused feature is also the size of $256\times 192$, and we use this fused feature to infer the missing part of the point cloud.
B. Comparison with Single View Reconstruction Method
Compared with the emergence of ViPC [56] in 2021, the current single view reconstruction has made a great breakthrough. To dispel the doubts about why we still need to study view guided point cloud completion instead of directly reconstructing shapes from images, we made a qualitative comparison with the latest single-view reconstruction model TripoSR [35] as fairness as possible. TripoSR is a 3D reconstruction model based on LRM (Large Reconstruction Model) [18] that can produce a 3D mesh from a single view image. Due to the need for a strong prior and smooth distribution of latent space for robust single view reconstruction, it can only restore a roughly reasonable shape but difficult to achieve fine surface reconstruction. In addition, single view reconstruction method relies heavily on the quality of the image. On the other hand, our method can achieve fine reconstruction of the missing parts based on the point cloud prior and supplemented by finely extracted image structure information. we uniformly sample 2048 points on the mesh surface and normalized the sampled point cloud into a unit sphere. Qualitative comparisons are shown in Fig. 9. It can be seen that TripoSR [35] tends to generate a shape with the outline of the image but lacks accurate depth on the image in the ShapeNet-ViPC dataset [56] in some categories. This indicates that it is unreliable to predict the complete shape based on the image alone. The point cloud completion task can use the image to infer the complete shape more reliably on the basis of the partial point cloud. Further more, relying on images to infer possible shapes is much more costly than point cloud completion. View guided point cloud completion [56] is a relatively cost-effective and reliable option.
Figure 9: Qualitative comparison with TripoSR [35]. It is difficult for single-view reconstruction to accurately estimate depth and predict the back in poor-quality images
C. Additional Qualitative Results
We provide additional completion results of our model in Fig. 10, Fig. 11, Fig. 12 and Fig. 13. All the point clouds in the figures are artificially rotated to an axis-aligned angle for easy observation. It can be seen that our method can give reasonable prediction in a variety of missing cases. Our model perform well on categories with large individual differences, such as lamps, chairs, etc. This is due to a deeper understanding of structural information, rather than learning the average shape distribution of the dataset.
Figure 10: Additional results on watercraft and table.
Figure 11: Additional results on sofa and lamp.
Figure 12: Additional results on car and plane.
Figure 13: Additional results on chair and cabinet.

### v3

@@ 2.1 Point cloud completion
The pioneer work of point cloud completion is PCN [51], which proposed a coarse-to-fine approach that is widely referenced in following studies [25, 21, 34, 28, 41, 56, 57, 43, 54, 36, 38, 37, 42, 33, 23, 4]. Though there are differences in feature extraction and utilization, the basic idea of these studies is to reconstruct the skeleton of the complete shape first and then refine it. PointTr [49] does not follow the coarse-to-fine manner but only generates the missing part of the partial point cloud. The idea of only generating the missing part also appeared in [48, 50]. In [32], a generative adversarial network is used for point cloud completion. PMP-Net [39] and PMP-Net++ [40] treat point cloud completion as a kind of deformation and complete point cloud by moving points to the right positions. P2C [5] introduces additional losses to supervise the latent expressions. Other unsupervised point cloud completion works [13, 52, 18, 3] achieve higher robustness through special training strategies Limited by inputs, these models must learn information about the complete shape from the occluded shape, which may lead to turning the task into a translation process from the occluded shape to the complete shape without meticulous analysis and design of the model.
@@ 4 Experimental Results
In this section, we first introduce the dataset and evaluation metrics in section 4.1. Quantitative and Qualitative comparisons are shown in section 4.2. Ablation studies are conducted in section 4.3. We also report the generalization ability of our method in section 4.4. Finally, the model complexity can be found in supplementary materials.
@@ 4.4.2 Results on Real Scenes
We also report qualitative results on KITTI [12] cars extracted by [51]. Results about qualitative results can be found in supplementary materials.